# PAT: Accelerating LLM Decoding via Prefix-Aware Attention with Resource Efficient Multi-Tile Kernel

### v2

@@ 5.2. Multi-tile Kernel
Figure 7. Multi-tile kernel design and validation on A100-SXM4-80GB GPU. (a) Global-to-shared memory transfer latency for varying data sizes (averaged over 1,000 reads). (b) Offline-selected tile-size configurations; check marks denote feasible settings, and circled numbers indicate violated constraints. (c) Average bandwidth utilization under different tile sizes. (d) Average kernel latency under different tile sizes.:
Insight and Approach. Given packed CTAs, PAT must choose per-CTA Q-tile $m$ and KV-tile $n$ to maximize resource efficiency. Because query sizes and KV lengths vary widely, single–tile-size kernels with padding (e.g., prior works (dao2023flashattention2; ye2025flashinfer; pan2025fasttree; zhu2024relayattention)) cannot adapt and suffer poor utilization (§3.3). To better exploit GPU shared memory and registers, we introduce a customized tiled attention kernel design that consists of: (1) a multi-tile kernel suite, where feasible $(m,n)$ configurations are derived from offline hardware and CTA-constraint analysis and implemented as resource-efficient kernels; and (2) a tile-size selector, an online decision-tree that selects the per-CTA $(m,n)$ to balance performance and parallelism.
@@ 1 $\mathcal{P}\leftarrow\emptyset$;
①Register and shared-memory constraints (upper bounds on $m,n$). Increasing $m$ or $n$ raises CTA shared-memory and register usage and can induce register spilling (nvidia2025cudaprogguide). To keep the kernel within hardware limits, we enforce two bounds: (1) Shared-memory constraint. One CTA’s shared-memory usage comprises the Q tile, K/V tile, and intermediate results (data type size $b^{\prime}$, usually higher precision). It must not exceed per-SM shared memory $S_{\mathrm{smem}}$:
@@
③CUTLASS constraint (lower bounds for $m$ and $n$): Efficient use of CUTLASS/CuTe MMA requires both tile sizes to be powers of two and at least [^5] It depends on data format; e.g., the minimum tile size is 32 for int8. 16 (cutlass_quickstart):
@@
Figure 8. Multi-tile kernel validation on H100-SXM4-80GB GPU. (a) Average bandwidth utilization under different tile sizes. (b) Average kernel latency under different tile sizes.:
@@
Figure 9. Execution pipeline of four CTAs under different concurrency. (a) High concurrency with dynamic KV lengths yields low tail efficiency and a large execution bubble $\mathbb{I}_{exe}$. (b–d) Lower concurrency cuts per-CTA latency and thus reduces $\mathbb{I}_{exe}$ across the shown execution orders.:
@@ 6. Multi Kernel Forward
After §5 packs the batch into CTAs and selects multi-tile kernels, we design two forward-stage strategies to mitigate kernel execution bubble $\mathbb{I}_{\mathrm{exe}}$3.3).
@@
Long KV Split. Multi-stream execution alone cannot eliminate execution bubbles in all cases, because some CTAs may have KV lengths that are orders of magnitude larger than others. We therefore adopt a KV-dimension splitting strategy similar to (pan2025fasttree). Specifically, we split any CTAs with KV length exceeding the mean KV length for all CTAs of the batch into equal parts to keep KV length under the mean value. This shortens the completion time of the last finishing CTAs and improves overall SM utilization of kernel execution.
@@ 7. Output Merge
We implement a lightweight merge kernel that uses online softmax (dao2022flashattention) to combine partial results at per-query granularity. Each CTA produces three per-query and per-head intermediates: a max score, a log-sum-exp accumulator, and a partial value-weighted sum. The merge kernel loads these intermediates from global memory, reduces them with online softmax, normalizes the accumulated value-weighted sum, concatenates all heads, and writes the final query output back to global memory. The small global read/write overhead for these intermediates is accounted for in the pack scheduler’s overhead analysis when deriving the end-to-end packing scheme (5.1).
@@ 8. Evaluation
Figure 10. Normalized kernel performance (higher is better) of PAT and the baselines for the attention computation across various decode batch configurations. The four panels from top to bottom show head configurations of 32/32, 16/8, 32/8, and 64/8. Missing bars arise because RelayAttention lacks support for multi-level or multiple first-level prefixes, and FastTree does not support the 16/8 and 64/8 head settings.:
@@
    What performance improvement does PAT achieve for online serving under real-world workloads? (§8.4)
@@ 8.1. Implementation of PAT
We implement PAT as a full attention kernel for the decode stage with about 3k lines of Cutlass/CuTe (nvidia2025cutlassdocs; nvidia2025cutequickstart) and C++ code. The multi-tile kernel (§5.2), multi-stream forward (§6), and merge kernel (§7) are built with Cutlass/CuTe. The asynchronous pack scheduler (§5.1) and API wrappers are implemented in C++. To overlap the data transfers with computation, all data movement from the global memory to the shared memory uses the cp_async primitive, together with double buffering (nvidia2025cudaprogguide). We expose the kernel API to Python via pybind11, and integrate it into vLLM (kwon2023pagedattention) (v0.9.0) as an off-the-shelf plugin with about 1.2k lines of Python code. PAT treats vLLM’s paged KV cache as its substrate: KV entries are managed as fixed-size blocks in block tables, and the pack scheduler operates only on block IDs produced by vLLM. This design lets PAT reuse vLLM’s existing KV paging implementation. To enable PAT in vLLM, only an environment variable VLLM_ATTENTION_BACKEND=PAT is required.
@@ 8.2. Experiment Setup
Models and Testbed. For end-to-end online serving (§8.4), we use two representative open-source models: Qwen3-8B (yang2025qwen3) and Llama3-8B (meta2024llama3). All experiments run on the NVIDIA A100-80GB-SXM4 GPU. Hardware info is summarized in Table 1. Software environment is CUDA 12.4 and PyTorch 2.7.0.
Baselines. We compare against five attention implementations spanning query-centric and KV-centric designs:
@@
    FlashAttention (dao2022flashattention; dao2023flashattention2) (v2.6.2): query-centric; maps each query to a CTA with a fixed tile size config $(64,128)$.
@@
    FlashInfer (ye2025flashinfer) (v0.2.5): query-centric; improves SM load balance by dynamic CTA partitioning; decoding tile config $(16,128)$.
@@
    FastTree (pan2025fasttree): KV-centric; uses a compute-oriented cost model to pack and reduce repeated KV loads; two tile configs $(64,32)$ and $(16,32)$.
@@
    RelayAttention (zhu2024relayattention): KV-centric; packs single-level shared prefixes across requests into CTAs to cut redundant memory accesses and runs them with FlashAttention’s kernel.
@@
Figure 11. End-to-end performance of PAT and the baselines under two models and two traces. Note that RelayAttention++ lacks support for multiple first-level prefixes, so its results on toolagent trace are unavailable. :
Kernel-Performance Workloads (§8.3). To compare kernel performance under identical batch and prefix structures, we construct synthetic decode batches as input, following (pan2025fasttree). Each decode batch is specified by $B$ and $L$. $B$ defines the prefix-tree structure and the number of leaves (i.e., batch size). e.g., $B=[1,4,16]$ yields two shared-prefix levels with 1 and 4 nodes, and 16 leaves. $L$ gives KV lengths per level. e.g., $L=[128,256,1024]$ sets level-1 and level-2 shared prefixes to 128 and 256 tokens, with 1024 non-shared tokens per request. We vary $(B,L)$ combinations to reflect different shared-prefix structures and batch settings as in Figure 10. Besides, we choose four head configurations (#heads, #kv_heads) common in Llama (meta2024llama3), Qwen (yang2025qwen3), and Gemma (google2025gemma) models: $(64,8)$, $(32,8)$, $(16,8)$, and $(32,32)$. The head dimension and data type are set to the commonly used 128 and FP16.
End-to-end workloads (§8.4 and §8.5). We evaluate PAT under online serving using vLLM (kwon2023pagedattention) (v0.9.0) on two real-world traces. (1) toolagent (qin2025mooncake): tool/agent workloads with task-specific system prompts (overall cache hit rate 59%). (2) conversation: combines the Meta-AI system instruction (total lengths 2517/2522 tokens for Llama3/Qwen3 tokenizers) with burstgpt prompts, following prior work (pan2025fasttree). We randomize language and country fields in the system instruction to create a three-level prefix whose lengths are 46, 348, 2123 (Llama3) or 45, 351, 2126 (Qwen3). We use only the first 30 minutes of each trace due to cost limit.
@@ 8.3. Kernel Performance
Overall results. Figure 10 reports normalized kernel performance (metric $\text{latency}^{-1}$, normalized to PAT) across decode-batch configurations. For configurations with shared prefixes ( 118), PAT achieves up to $21.5\times$, $11.7\times$, $3.2\times$, $11.9\times$ and $5.7\times$ speedups over FlashAttention, FlashInfer, FastTree, RelayAttention and RelayAttention++, respectively, across four attention-head settings. These gains arise from three factors: (1) a prefix-aware packing scheduler that cuts redundant global KV accesses (§5.1); (2) a multi-tile kernel combined with an online tile-size selector that adapts $(m,n)$ per CTA to better use bandwidth and on-chip memory (§5.2); and (3) multi-stream forward execution and long-KV splitting that mitigate execution bubbles caused by multi-tile kernels and KV-length dynamics (§6). Together, these designs make PAT consistently more efficient than the baselines.
@@
Compared with KV-centric kernels. FastTree, which also targets multi-level shared prefixes, remains the strongest baseline but is still $3.8\%$$68.9\%$ slower than PAT for two reasons: (1) its compute-oriented packing cost model is ill-suited to memory-bound decode attention (§2.3 and §3.2); and (2) its double-tile approach launches two kernels serially as in Figure 13(b), introducing execution bubbles. By contrast, PAT uses a memory-oriented packing strategy plus multi-stream forward to avoid these inefficiencies. RelayAttention++ cuts latency by $67.4\%$ versus RelayAttention, confirming that L2 plus KV-cache reuse reduces redundant global loads, yet RelayAttention++ is still $1.7\times$ slower than PAT, showing that L2 cache alone cannot fully eliminate redundant KV loads.
@@ 8.4. End-to-End Comparison
Why baselines fail. PAT achieves significant gains for its prefix-award pack scheduler, multi-tile kernel, and multi-stream forward. In contrast, RelayAttention++ supports only a single-level system prefix and delegates the forward kernel to FlashAttention, so its curves largely track FlashAttention, which can not mitigate redundant KV loads at all. Model configuration also matters. Llama-3-8B’s 8K context length limit (vs. 32K for Qwen3-8B) requires us to filter ultra-long requests, which reduces the requests with long non-prefix context and leads to lower absolute latency for both RelayAttention++ and PAT on Llama-3-8B compared with that of Qwen3-8B. Besides, FlashInfer improves SM utilization via long-CTA-splitting load balance, which helps at low request rate while adding scheduling overhead that grows with request rate, resulting in the highest TPOT.
@@ 8.5. Ablation Study
Figure 12. Results of ablation baselines. (a) Average attention latency. (b) Global memory read and write size.:
Figure 13. CTA execution pipeline on $SM_{0}-SM_{5}$ with two tile size configurations (collected by PTX (nvidia2023ptxisa)), where white space represents execution bubbles. (a) Multi-stream execution pipeline. (b) Serial execution pipeline.:
To evaluate the contribution of each design in PAT, we build the following ablation baselines: (1) PAT-compute, which adopts the cost model from FastTree (pan2025fasttree) for the packing scheduler; (2) PAT-naive, which simply packs each node in the tree-structured block table into a CTA; (3) PAT-fixed, which disables the multi-tile kernel of PAT and instead uses the fixed tile configuration $(64,128)$ as in FlashAttention (shah2024flashattention3); (4) PAT-serial, which disables the multi-stream forward mechanism of PAT and adopts the serial multi-kernel execution similar to FastTree (pan2025fasttree). We use the same synthetic traces as in §8.3 and adopt the attention head configuration of Llama-3-8B to compare PAT with ablation baselines.
@@
Importance of multi-tile kernel. As shown in Figure 12(a), enforcing a fixed tile size (PAT-fixed) increases attention latency by 39% compared to PAT. In the ablation workload, CTAs exhibit query sizes from 1 to 64 and KV lengths from 32 to 4096, which makes one-size-fits-all kernels highly inefficient. Prior kernels (shah2024flashattention3; ye2025flashinfer; zhu2024relayattention; yao2024deft) adopt such fixed designs, leading to padding overhead and execution bubbles. In contrast, PAT’s multi-tile kernel enables efficient adaptation to diverse CTA configurations.
Effectiveness of multi-stream forward. As shown in Figure 12(a), the average attention latency of PAT-serial is 4.8% higher than PAT. This is because serial execution aggravates execution bubbles. For example, in a two-level prefix decode batch (Figure 10 6), Figure 13(a) and Figure 13(b) show the CTA execution pipelines of PAT and PAT-serial, where PAT-serial suffers from substantial memory waste and execution bubbles. By running multi-tile kernels across multiple CUDA streams in parallel, the multi-stream forward strategy of PAT effectively mitigates execution bubbles caused by workload dynamicity and achieves lower attention latency.

### v3

@@ 5.2. Multi-tile Kernel
Figure 8. Multi-tile kernel design and validation on A100-SXM4-80GB GPU. (a) Global-to-shared memory transfer latency for varying data sizes (averaged over 1,000 reads). (b) Offline-selected tile-size configurations; check marks denote feasible settings, and circled numbers indicate violated constraints. (c) Average bandwidth utilization under different tile sizes. (d) Average kernel latency under different tile sizes.
Insight and Approach. Given packed CTAs, PAT must choose per-CTA Q-tile $m$ and KV-tile $n$ to maximize resource efficiency. Because query sizes and KV lengths vary widely, single–tile-size kernels with padding (e.g., prior works (Dao, 2023; Ye et al., 2025; Pan et al., 2025; Zhu et al., 2024)) cannot adapt and suffer poor utilization (§3.3). To better exploit GPU shared memory and registers, we introduce a customized tiled attention kernel design that consists of: (1) a multi-tile kernel suite, where feasible $(m,n)$ configurations are derived from offline hardware and CTA-constraint analysis and implemented as resource-efficient kernels; and (2) a tile-size selector, an online decision-tree that selects the per-CTA $(m,n)$ to balance performance and parallelism.
@@ 1 $\mathcal{P}\leftarrow\emptyset$;
①Register and shared-memory constraints (upper bounds on $m,n$). Increasing $m$ or $n$ raises CTA shared-memory and register usage and can induce register spilling (NVIDIA, 2025b). To keep the kernel within hardware limits, we enforce two bounds: (1) Shared-memory constraint. One CTA’s shared-memory usage comprises the Q tile, K/V tile, and intermediate results (data type size $b^{\prime}$, usually higher precision). It must not exceed per-SM shared memory $S_{\mathrm{smem}}$:
@@
③CUTLASS constraint (lower bounds for $m$ and $n$): Efficient use of CUTLASS/CuTe MMA requires both tile sizes to be powers of two and at least [^5] It depends on data format; e.g., the minimum tile size is 32 for int8. 16 (NVIDIA, 2025c):
@@
Figure 9. Multi-tile kernel validation on H100-SXM4-80GB GPU. (a) Average bandwidth utilization under different tile sizes. (b) Average kernel latency under different tile sizes.
@@
Figure 10. Execution pipeline of four CTAs under different concurrency. (a) High concurrency with dynamic KV lengths yields low tail efficiency and a large execution bubble $\mathbb{I}_{exe}$. (b–d) Lower concurrency cuts per-CTA latency and thus reduces $\mathbb{I}_{exe}$ across the shown execution orders.
@@ 6. Multi Kernel Forward
After §5 packs the batch into CTAs and selects multi-tile kernels, we design two forward-stage strategies to mitigate kernel execution bubbles $\mathbb{I}_{\mathrm{exe}}$3.3).
@@
Long KV Split. Multi-stream execution alone cannot eliminate execution bubbles in all cases, because some CTAs may have KV lengths that are orders of magnitude larger than others. We therefore adopt a KV-dimension splitting strategy similar to (Pan et al., 2025). Specifically, we split any CTAs with KV length exceeding the mean KV length for all CTAs of the batch into equal parts to keep the KV length under the mean value. This shortens the completion time of the last finishing CTAs and improves overall SM utilization of kernel execution.
@@ 7. Output Merge
We implement a lightweight merge kernel that uses online softmax (Dao et al., 2022) to combine partial results at per-query granularity. Each CTA produces three per-query and per-head intermediates: a max score, a log-sum-exp accumulator, and a partial value-weighted sum. The merge kernel loads these intermediates from global memory, reduces them with online softmax, normalizes the accumulated value-weighted sum, concatenates all heads, and writes the final query output back to global memory. The small global read/write overhead for these intermediates is accounted for in the pack scheduler’s overhead analysis when deriving the end-to-end packing scheme (5.1).
@@ 8. Evaluation
Figure 11. Normalized kernel performance (higher is better) of PAT and the baselines for the attention computation across various decode batch configurations on NVIDIA A100 GPU (80GB). The four panels from top to bottom show head configurations (num_attention_heads/num_key_value_headsnum\_attention\_heads/num\_key\_value\_heads) of 32/32, 16/8, 32/8, and 64/8. Missing bars arise because RelayAttention lacks support for multi-level or multiple first-level prefixes, and FastTree does not support the 16/8 and 64/8 head settings.
@@
    What performance improvement does PAT achieve for online serving under real-world workloads? (§8.4, §8.5)
@@ 8.1. Implementation of PAT
We implement PAT as a full attention kernel for the decode stage with about 3k lines of Cutlass/CuTe (NVIDIA, 2025f, e) and C++ code. The multi-tile kernel (§5.2), multi-stream forward (§6), and merge kernel (§7) are built with Cutlass/CuTe. The asynchronous pack scheduler (§5.1) and API wrappers are implemented in C++. To overlap the data transfers with computation, all data movement from the global memory to the shared memory uses the cp_async primitive, together with double buffering (NVIDIA, 2025b). We expose the kernel API to Python via pybind11, and integrate it into vLLM (Kwon et al., 2023) (v0.9.0) as an off-the-shelf plugin with about 1.2k lines of Python code. PAT treats vLLM’s paged KV cache as its substrate: KV entries are managed as fixed-size blocks in block tables, and the pack scheduler operates only on block IDs produced by vLLM. This design lets PAT reuse vLLM’s existing KV paging implementation. To enable PAT in vLLM, only an environment variable VLLM_ATTENTION_BACKEND=PAT is required.
@@ 8.2. Experiment Setup
Models and Testbed. For kernel benchmark (§8.3), we evaluate PAT on both NVIDIA A100 GPU (80GB) and NVIDIA H100 GPU. For end-to-end online serving (§8.4), we use two representative LLMs, Qwen3-8B and Llama-3-8B, on a single A100 GPU. We further evaluate PAT under distributed settings and Mixture-of-Experts (MoE) architectures in §8.5 using Qwen2.5-72B-Instruct on four A100 GPUs and Qwen3-30B-A3B on one A100 GPU. The software environment is CUDA 12.4 and PyTorch 2.7.0.
Baselines. We compare against seven attention implementations spanning query-centric and KV-centric designs:
@@
    FlashAttention (Dao et al., 2022; Dao, 2023) (v2.5.9): query-centric; maps each query to a CTA with a fixed tile size config $(64,128)$.
@@
    FlashInfer (Ye et al., 2025) (v0.2.5): query-centric; improves SM load balance by dynamic CTA partitioning; decoding tile config $(16,128)$.
@@
    FastTree (Pan et al., 2025): KV-centric; uses a compute-oriented cost model to pack and reduce repeated KV loads; two tile configs $(64,32)$ and $(16,32)$.
@@
    RelayAttention (Zhu et al., 2024): KV-centric; packs first-level shared prefixes into CTAs to cut redundant memory accesses and runs them with FlashAttention’s kernel.
@@
  6. (6)
    DeFT (Yao et al., 2024): KV-centric, aggregates queries with shared KV and adjusts the KV length in each CTA for load balance; fixed tile size config $(32,16)$.
  7. (7)
    Cascade Inference (Ye et al., 2024a): KV-centric, packs prefixes into CTAs using fixed settings.
Kernel-Performance Workloads (§8.3). To compare kernel performance under identical batch and prefix structures, we construct synthetic decode batches as input, following (Pan et al., 2025). Each decode batch is specified by $B$ and $L$. $B$ defines the prefix-tree structure and the number of leaves (i.e., batch size). e.g., $B=[1,4,16]$ yields two shared-prefix levels with 1 and 4 nodes, and 16 leaves. $L$ gives KV lengths per level. e.g., $L=[128,256,1024]$ sets level-1 and level-2 shared prefixes to 128 and 256 tokens, with 1024 non-shared tokens per request. We vary $(B,L)$ combinations to reflect different shared-prefix structures and batch settings as in Figure 11. Besides, we choose four head configurations (#heads, #kv_heads) common in Llama (AI@Meta, 2024), Qwen (Yang et al., 2025), and Gemma (Team, 2025) models: $(64,8)$, $(32,8)$, $(16,8)$, and $(32,32)$. The head dimension and data type are set to the commonly used 128 and FP16.
End-to-end workloads (§8.4 and §8.6). We evaluate PAT under online serving using vLLM (Kwon et al., 2023) (v0.9.0) on two real-world traces. (1) toolagent (Qin et al., 2025): tool/agent workloads with task-specific system prompts (overall cache hit rate 59%). (2) conversation: combines the Meta-AI system instruction (total lengths 2517/2522 tokens for Llama3/Qwen3 tokenizers) with burstgpt prompts, following prior work (Pan et al., 2025). We randomize language and country fields in the system instruction to create a three-level prefix whose lengths are 46, 348, 2123 (Llama3) or 45, 351, 2126 (Qwen3). We use only the first 30 minutes of each trace due to cost limit.
Figure 12. End-to-end performance of PAT and the baselines under two models and two traces. Note that RelayAttention++ lacks support for multiple first-level prefixes, so its results on toolagent trace are unavailable.
@@ 8.3. Kernel Performance
Overall results. Figure 11 reports normalized kernel performance (metric $\text{latency}^{-1}$, normalized to PAT) across decode-batch configurations on NVIDIA A100 GPU (80GB) [^7] Evaluation results on NVIDIA H100 GPU are presented in Appendix A.. For configurations with shared prefixes ( 118), PAT achieves up to $21.5\times$, $11.7\times$, $3.2\times$, $11.9\times$ and $5.7\times$ speedups over FlashAttention, FlashInfer, FastTree, RelayAttention and RelayAttention++, respectively, across four attention-head settings. These gains arise from three factors: (1) a prefix-aware packing scheduler that cuts redundant global KV accesses (§5.1); (2) a multi-tile kernel combined with an online tile-size selector that adapts $(m,n)$ per CTA to better use bandwidth and on-chip memory (§5.2); and (3) multi-stream forward execution and long-KV splitting that mitigate execution bubbles caused by multi-tile kernels and KV-length dynamics (§6). Together, these designs make PAT consistently more efficient than the baselines.
@@
Compared with KV-centric kernels. FastTree, which also targets multi-level shared prefixes, remains the strongest baseline but is still $3.8\%$$68.9\%$ slower than PAT for two reasons: (1) its compute-oriented packing cost model is ill-suited to memory-bound decode attention (§2.3 and §3.2); and (2) its double-tile approach launches two kernels serially as in Figure 15(b), introducing execution bubbles. By contrast, PAT uses a memory-oriented packing strategy plus multi-stream forward to avoid these inefficiencies. RelayAttention++ cuts latency by $67.4\%$ versus RelayAttention, confirming that L2 plus KV-cache reuse reduces redundant global loads, yet RelayAttention++ is still $1.7\times$ slower than PAT, showing that L2 cache alone cannot fully eliminate redundant KV loads. In addition, DeFT and Cascade Inference both use a naive packing scheme for shared prefixes. While DeFT’s load-balancing reduces SM execution bubbles from long-tail CTAs, neither method effectively reduces global memory accesses, which dominate decode attention latency. Therefore, PAT achieves $76.6\%$ and $41.2\%$ lower attention latency on average than DeFT and Cascade Inference, respectively.
@@ 8.4. End-to-End Comparison
Why baselines fail. PAT achieves significant gains for its prefix-aware pack scheduler, multi-tile kernel, and multi-stream forward. In contrast, RelayAttention++ supports only a single-level system prefix and delegates the forward kernel to FlashAttention, so its curves largely track FlashAttention, which can not mitigate redundant KV loads at all. Model configuration also matters. Llama-3-8B’s 8K context length limit (vs. 32K for Qwen3-8B) requires us to filter ultra-long requests, which reduces the requests with long non-prefix context and leads to lower absolute latency for both RelayAttention++ and PAT on Llama-3-8B compared with that of Qwen3-8B. Besides, FlashInfer improves SM utilization via long-CTA-splitting load balance, which helps at low request rate while adding scheduling overhead that grows with request rate, resulting in the highest TPOT.
8.5. Distributed and MoE Extensions
Figure 13. End-to-end performance of PAT and baselines under TP/$P$ and MoE architectures.
Recent production LLM deployments increasingly rely on multi-GPU inference as model weights exceed a single GPU’s memory capacity, and many models adopt Mixture-of-Experts (MoE) architectures to increase capacity at moderate cost. PAT seamlessly supports tensor parallelism (TP), pipeline parallelism (PP), and MoE: TP and PP partition attention heads or transformer blocks across GPUs, while each device retains full KV cache for its assigned heads or blocks; MoE architecture replaces the FFN layer with multiple experts but leaves attention execution unchanged. Therefore, PAT’s pack scheduler and multi-tile kernel run unchanged.
To validate these properties, we evaluate PAT under the toolagent trace using Qwen2.5-72B-Instruct with TP=2 and PP=2 on four A100 GPUs, and Qwen3-30B-A3B on one A100 GPU. As shown in Figure 13, compared to three baselines, PAT reduces average TPOT by $14.3-26.7\%$ on Qwen2.5-72B-Instruct and by $5.53-16.9\%$ on Qwen3-30B-A3B. These results suggest that prefix-aware packing and the multi-tile kernel remain effective under common distributed and MoE deployment configurations.
@@ 8.6. Ablation Study
Figure 14. Results of ablation baselines. (a) Average attention latency. (b) Global memory read and write size.
Figure 15. CTA execution pipeline on $SM_{0}-SM_{5}$ with two tile size configurations (collected by PTX (NVIDIA, 2023)), where white space represents execution bubbles. (a) Multi-stream execution pipeline. (b) Serial execution pipeline.
To evaluate the contribution of each design in PAT, we build the following ablation baselines: (1) PAT-compute, which adopts the cost model from FastTree (Pan et al., 2025) for the packing scheduler; (2) PAT-naive, which simply packs each node in the tree-structured block table into a CTA; (3) PAT-fixed, which disables the multi-tile kernel of PAT and instead uses the fixed tile configuration $(64,128)$ as in FlashAttention (Shah et al., 2024); (4) PAT-serial, which disables the multi-stream forward mechanism of PAT and adopts the serial multi-kernel execution similar to FastTree (Pan et al., 2025). We use the same synthetic traces as in §8.3 and adopt the attention head configuration of Llama-3-8B to compare PAT with ablation baselines.
@@
Importance of multi-tile kernel. As shown in Figure 14(a), enforcing a fixed tile size (PAT-fixed) increases attention latency by 39% compared to PAT. In the ablation workload, CTAs exhibit query sizes from 1 to 64 and KV lengths from 32 to 4096, which makes one-size-fits-all kernels highly inefficient. Prior kernels (Shah et al., 2024; Ye et al., 2025; Zhu et al., 2024; Yao et al., 2024) adopt such fixed designs, leading to padding overhead and execution bubbles. In contrast, PAT’s multi-tile kernel enables efficient adaptation to diverse CTA configurations.
Effectiveness of multi-stream forward. As shown in Figure 14(a), the average attention latency of PAT-serial is 4.8% higher than PAT. This is because serial execution aggravates execution bubbles. For example, in a two-level prefix decode batch (Figure 11 6), Figure 15(a) and Figure 15(b) show the CTA execution pipelines of PAT and PAT-serial, where PAT-serial suffers from substantial memory waste and execution bubbles. At runtime, PAT’s multi-tile kernel selects a tile configuration for each CTA from the equivalent configuration set. In our A100 experiments, each batch typically uses 1–5 of the 11 available configurations (Figure 8). PAT’s multi-stream forward then launches a separate CUDA stream for CTAs with each active tile configuration, allowing these streams to run in parallel under the GPU scheduler. With multiple configurations, some launch bubbles remain, as indicated by the blank region on the left of Figure 15(a). Nevertheless, parallel streams substantially reduce execution bubbles compared to serial execution, thereby lowering attention latency.